5. A sample of gas has a volume of 215 cm3 at 23.5 °C and 84.6 kPa. What volume will the gas occupy at STP? to 192
Answer #1

V = 215cm^3 = 215ml = 0.215L

T   = 23.5+273   = 296.5K

P = 84.6KPa

experimental conditions           STP conditions

V1 = 0.215L    V2 =

T1 = 296.5K T2 = 273K

P1 = 84.6KPa    P2 = 101.325KPa

   P1V1/T1    = P2V2/T2

V2    = P1V1T2/T1P2

= 84.6*0.215*273/(296.5*101.325)

                                  = 0.615L

volume of gas STP   = 0.615L = 615cm^3